# Derivative-Free Guidance in Continuous and Discrete Diffusion Models with Soft Value-Based Decoding

### v3

@@ Abstract
Diffusion models excel at capturing the natural design spaces of images, molecules, DNA, RNA, and protein sequences. However, rather than merely -generating designs that are natural, we often aim to optimize downstream reward functions while preserving the naturalness of these design spaces. Existing methods for achieving this goal often require “differentiable” proxy models (e.g., classifier guidance or DPS) or involve computationally expensive fine-tuning of diffusion models (e.g., classifier-free guidance, RL-based fine-tuning). In our work, we propose a new method to address these challenges. Our algorithm is an iterative sampling method that integrates soft value functions, which looks ahead to how intermediate noisy states lead to high rewards in the future, into the standard inference procedure of pre-trained diffusion models. Notably, our approach avoids fine-tuning generative models and eliminates the need to construct differentiable models. This enables us to (1) directly utilize non-differentiable features/reward feedback, commonly used in many scientific domains, and (2) apply our method to recent discrete diffusion models in a principled way. Finally, we demonstrate the effectiveness of our algorithm across several domains, including image generation, molecule generation, and DNA/RNA sequence generation. The code is available at https://github.com/masa-ue/SVDD.
@@ 1 Introduction
Diffusion models have gained popularity as powerful generative models. Their applications extend beyond image generation to include natural language generation (Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2023), molecule generation (Jo et al., 2022; Vignac et al., 2022), and biological (DNA, RNA, protein) sequence generation (Avdeyev et al., 2023; Stark et al., 2024), and in each of these domains diffusion models have been shown to be very effective at capturing complex natural distributions. However, in practice, we might not only want to generate realistic samples, but to produce samples that optimize specific downstream reward functions while preserving naturalness by leveraging pre-trained models. For example, in computer vision, we might aim to generate natural images with high aesthetic and alignment scores. In drug discovery, we may seek to generate valid molecules with high QED/SA/docking scores (Lee et al., 2023; Jin et al., 2018) or natural RNAs (such as mRNA vaccines (Cheng et al., 2023)) with high translational efficiency and stability (Castillo-Hair and Seelig, 2021; Asrani et al., 2018), and natural DNAs with high cell-specificity (Gosai et al., 2023; Taskiran et al., 2024; Lal et al., 2024).
@@
The optimization of downstream reward functions using pre-trained diffusion models has been approached in various ways. In our work, we focus on non-fine-tuning-based methods because fine-tuning generative models (e.g., when using classifier-free guidance (Ho et al., 2020) or RL-based fine-tuning (Black et al., 2023; Fan et al., 2023; Uehara et al., 2024; Clark et al., 2023; Prabhudesai et al., 2023)) often becomes computationally intensive, especially as pre-trained generative models grow larger in the era of “foundation models”. Although classifier guidance and its variants (e.g., Dhariwal and Nichol (2021); Song et al. (2020); Chung et al. (2022); Bansal et al. (2023); Ho et al. (2022)) have shown success as non-fine-tuning methods in these settings, they face significant challenges. Firstly, as they would require constructing differentiable proxy models, they cannot directly incorporate useful domain-specific non-differentiable features (e.g., molecular/protein descriptors (van Westen et al., 2013; Ghiringhelli et al., 2015; Gainza et al., 2020)) or non-differentiable reward feedback (e.g., physics-based simulations such as Vina and Rosetta (Trott and Olson, 2010; Alhossary et al., 2015; Alford et al., 2017)). This limitation also hinders the application of current classifier guidance methods to recently developed discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2023) in a principled manner (i.e., without transforming the discrete space into the Euclidean space). Additionally, gradient computation can be memory-intensive with large proxy models.
@@
Figure 1: Summary of SVDD. The notation $v$ denotes value functions that predict reward $r(x_{0})$ (at time 00) from states at time $t-1$. SVDD involves two steps: (1) generating multiple noisy states from pre-trained models and (2) selecting the state with the highest value according to the value function.
@@
We propose a novel method, SVDD (Soft Value-based Decoding in Diffusion models), for optimizing downstream reward functions in diffusion models (summarized in Figure 1) to address the aforementioned challenges. Inspired by recent literature on RL-based fine-tuning (Uehara et al., 2024), we first introduce soft value functions that serve as look-ahead functions, indicating how intermediate noisy samples lead to high rewards in the future of the diffusion denoising process. After learning (or approximating) these value functions, we present a new inference-time technique, SVDD, which obtains multiple noisy states from the policy (i.e., denoising map) of pre-trained diffusion models and selects the sample with the highest value function at each time step. Specifically, we introduce two algorithms (SVDD-MC and SVDD-PM) depending on how we estimate value functions. Notably, the SVDD-PM approach, leveraging the characteristics of diffusion models (e.g., utilizing the forward process in diffusion models to directly map $t$ to 00 in terms of expectation in Figure 1), requires no additional training as long as we have access to the reward feedback.
@@
Table 1: A comparison of our method with a number of prior approaches. “Non-differentiable” refers to the method’s capability to operate without requiring differentiable proxy models. “No Training” means that no additional training of the diffusion model is required as long as we have access to the reward feedback. Regarding the comparison with standard SMC-based methods, refer to Section A.
@@
Classifier guidance (Dhariwal and Nichol, 2021) $\checkmark$
@@
Classifier-free (Ho and Salimans, 2022) $\checkmark$
@@
SVDD-MC (Ours) $\checkmark$ $\checkmark$
SVDD-PM (Ours) $\checkmark$ $\checkmark$ $\checkmark$
Our contributions are summarized as follows (also see Table 1). We propose a novel technique for optimizing downstream reward functions in pre-trained diffusion models that eliminates the need to construct differentiable proxy models and avoids the need to fine-tune the generative model itself. The first property allows for the use of non-differentiable reward feedback, which is common in many scientific fields, and makes our method applicable to recent discrete diffusion models without any modification. The second property addresses the high computational cost associated with fine-tuning diffusion models. We demonstrate the effectiveness of our methods across various domains, including image generation, molecule generation, and DNA/RNA generation.
@@ 2 Related Works
Here, we summarize related work. We first outline methods relevant to our goal, categorizing them based on whether they involve fine-tuning. We then discuss related directions, such as discrete diffusion models, where our method excels, and decoding in autoregressive models.
@@
There are primarily two methods for optimizing downstream functions in diffusion models without fine-tuning. Regarding the comparison with the most related works based on sequential Monte Carlo (SMC), we defer the discussion to Section 5.4 and Section A.
  • Classifier guidance (Dhariwal and Nichol, 2021; Song et al., 2020): It has been widely used to condition pre-trained diffusion models without fine-tuning. Although these methods do not originally focus on optimizing reward functions, they can be applied for this purpose (Uehara et al., 2024, Section 6.2). In this approach, an additional derivative of a certain value function is incorporated into the drift term (mean) of pre-trained diffusion models during inference. Subsequent variants (e.g., Chung et al. (2022); Ho et al. (2022); Bansal et al. (2023); Guo et al. (2024); Wang et al. (2022); Yu et al. (2023)) have been proposed to simplify the learning of value functions. However, these methods require constructing differentiable models, which limits their applicability to non-differentiable features/reward feedbacks commonly encountered in scientific domains as mentioned in Section 1. Additionally, this approach cannot be directly extended to discrete diffusion models in a principle way  [^1] We note that a notable exception has been recently proposed (Nisonoff et al., 2024). However, our method can be applied to both continuous and discrete diffusion models in a unified manner. Furthermore, their method in practice requires the differentiability of proxy models, unlike our proposal. Regarding more details, refer to Section B. . Our approach aims to address these challenges.
  • Best-of-N: The naive way is to generate multiple samples and select the top samples based on the reward functions, known as best-of-N in the literature on (autoregressive) LLMs (Stiennon et al., 2020; Nakano et al., 2021; Touvron et al., 2023; Beirami et al., 2024; Gao et al., 2023). This approach is significantly less efficient than ours, as our method uses soft-value functions that predict how intermediate noisy samples lead to high rewards in the future. We validate this experimentally in Section 6.
@@
Several methods exist for fine-tuning generative models to optimize downstream reward functions, such as classifier-free guidance (Ho and Salimans, 2022) and RL-based fine-tuning (Fan et al., 2023; Black et al., 2023)/its variants (Dong et al., 2023). However, these approaches often come with caveats, including high computational costs and the risk of easily forgetting pre-trained models. In our work, we propose an inference-time technique that eliminates the need for fine-tuning generative models, similar to classifier guidance.
@@
Based on seminal works Austin et al. (2021); Campbell et al. (2022), recent work on masked diffusion models (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024) has demonstrated their strong performance in natural language generation. Additionally, they have been applied to biological sequence generation (e.g., DNA, protein sequences in Campbell et al. (2024); Sarkar et al. (2024)). In these cases, the use of diffusion models over autoregressive models is particularly apt, given that many biological sequences ultimately adopt complex three-dimensional structures. We also note that ESM3 (Hayes et al., 2024), a widely recognized foundational model in protein sequence generation, bears similarities to masked diffusion models.
Despite its significance, it cannot be integrated with standard classifier guidance because adding a continuous gradient to a discrete objective is not inherently valid. Unlike standard classifier guidance, our algorithm can be seamlessly applied to discrete diffusion models.
Decoding in autoregressive models with rewards.
@@ 3 Preliminaries and Goal
In this section, we describe the standard method for training diffusion models and outline the objective of our work: optimizing downstream reward functions given pre-trained diffusion models.
@@ 3.1 Diffusion Models
In diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020), our goal is to learn a sampler $p(x)\in\Delta(\mathcal{X})$ given data consisting of $(x)\in\mathcal{X}$. The training process for a standard diffusion model is summarized as follows. First, we introduce a (fixed) forward process $q_{t}:\mathcal{X}\to\Delta(\mathcal{X})$. Now, after introducing the forward process, we aim to learn a backward process: $\{p_{t}\}$ where each $p_{t}$ is $\mathcal{X}\to\Delta(\mathcal{X})$ so that the induced distributions induced by the forward process and backward process match marginally. For this purpose, by parametrizing the backward processes with $\theta\in\mathbb{R}^{d}$, we typically use the following loss function:
@@
which is derived from the variational lower bound of the negative likelihood (i.e., ELBO).
@@
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. Here, $\hat{x}_{0}(x_{t};\theta)$ is a neural network that predicts $x_{0}$ from $x_{t}$ (i.e., $\mathbb{E}_{q}[x_{0}|x_{t}]$).
@@ 3.2 Objective: Generating Samples with High Rewards While Preserving Naturalness
We consider a scenario where we have a pre-trained diffusion model, which is trained using the loss function explained in Section 3.1. These pre-trained models are typically designed to excel at characterizing the natural design space (e.g., image space, biological space, or chemical space) by emulating the extensive training dataset. Our work focuses on obtaining samples that also optimize downstream reward functions $r:\mathcal{X}\to\mathbb{R}$ (e.g., QED and SA in molecule generation) while maintaining the naturalness by leveraging pre-trained diffusion models. We formalize this goal as follows.
@@
$p^{(\alpha)}(x)$ $:=\mathop{\mathrm{argmax}}_{p\in[\Delta(\mathcal{X})]}\underbrace{\mathbb{E}_{x\sim p(\cdot)}[r(x)]}_{\textbf{term (a)}}-\alpha\underbrace{\mathrm{KL}(p(\cdot)\|p^{\mathrm{pre}}(\cdot))}_{\textbf{term(b)}}$
$\propto\exp(r(x)/\alpha)p^{\mathrm{pre}}(x).$
Here, the term (a) is introduced to optimize reward function, while the term (b) is used to maintain the naturalness of the generated samples.
@@
Several existing approaches target this goal (or its variant), as discussed in Section 2, including classifier guidance, fine-tuning (RL-based or classifier-free), and Best-of-N. In our work, we focus on non-fine-tuning-based methods, such as classifier guidance and Best-of-N. Specifically, we aim to address the limitations of these methods: the requirement for differentiable proxy models in classifier guidance and the inefficiency of Best-of-N.
Remark 1.
We can easily extend all the diffusion models discussed in our paper to cases where the pre-trained model is a conditional diffusion model.

### v4

@@ Abstract
Diffusion models excel at capturing the natural design spaces of images, molecules, and biological sequences of DNA, RNA, and proteins. However, for many applications from biological research to biotherapeutic discovery, rather than merely generating designs that are natural, we aim to optimize downstream reward functions while preserving the naturalness of these design spaces. Existing methods for achieving this goal often require “differentiable” proxy models (e.g., classifier guidance) or computationally-expensive fine-tuning of diffusion models (e.g., classifier-free guidance, RL-based fine-tuning). Here, we propose a new method, Soft Value-based Decoding in Diffusion models (SVDD), to address these challenges. SVDD is an iterative sampling method that integrates soft value functions, which looks ahead to how intermediate noisy states lead to high rewards in the future, into the standard inference procedure of pre-trained diffusion models. Notably, SVDD avoids fine-tuning generative models and eliminates the need to construct differentiable models. This enables us to (1) directly utilize non-differentiable features/reward feedback, commonly used in many scientific domains, and (2) apply our method to recent discrete diffusion models in a principled way. Finally, we demonstrate the effectiveness of SVDD across several domains, including image generation, molecule generation (optimization of docking scores, QED, SA), and DNA/RNA generation (optimization of activity levels). The code is available at https://github.com/masa-ue/SVDD.
@@ 1 Introduction
Diffusion models have gained popularity as powerful generative models. Their applications extend beyond image generation to include natural language generation (Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2023), molecule generation (Jo et al., 2022; Vignac et al., 2022), and biological (DNA, RNA, protein) sequence generation (Avdeyev et al., 2023; Stark et al., 2024). In each of these domains, diffusion models have been shown to be very effective at capturing complex natural distributions. However, in practice, we might not only want to generate realistic samples, but to produce samples that optimize specific downstream reward functions while preserving naturalness by leveraging pre-trained models. For example, in computer vision, we might aim to generate natural images with high aesthetic and alignment scores (Black et al., 2023; Fan et al., 2023). In drug discovery, we may seek to generate valid molecules with high QED/SA/docking scores (Lee et al., 2023; Jin et al., 2018) or RNAs (such as mRNA vaccines (Cheng et al., 2023)) with high translational efficiency and stability (Castillo-Hair and Seelig, 2021; Asrani et al., 2018), and regulatory DNAs that drives high cell-specificity of expression (Gosai et al., 2023; Taskiran et al., 2024; Lal et al., 2024).
@@
The optimization of downstream reward functions using pre-trained diffusion models has been approached in various ways. In our work, we focus on non-fine-tuning-based methods because fine-tuning generative models (e.g., when using classifier-free guidance (Ho et al., 2020) or RL-based fine-tuning (Black et al., 2023; Fan et al., 2023; Uehara et al., 2024; Clark et al., 2023; Prabhudesai et al., 2023)) often becomes computationally intensive, especially as pre-trained generative models grow larger in the era of “foundation models”. Although classifier guidance and its variants (e.g., Dhariwal and Nichol (2021); Song et al. (2020); Chung et al. (2022); Bansal et al. (2023); Ho et al. (2022)) have shown some success as non-fine-tuning methods in these settings, they face significant challenges. First, as they would require constructing differentiable proxy models, they cannot directly incorporate useful non-differentiable features (e.g., molecular/protein descriptors (van Westen et al., 2013; Ghiringhelli et al., 2015; Gainza et al., 2020)) or non-differentiable reward feedback (e.g., physics-based simulations such as Vina and Rosetta (Trott and Olson, 2010; Alhossary et al., 2015; Alford et al., 2017)), which are particularly important in molecule design to optimize docking scores, stability, etc. This limitation also hinders the principled application of current classifier guidance methods to recently-developed discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Lou et al., 2023) (i.e., without transforming the discrete space into the Euclidean space).
@@
Figure 1: Summary of SVDD. $v$ denotes value functions that predict reward $r(x_{0})$ (at time 00) from states at time $t-1$. SVDD involves two steps: (1) generating multiple noisy states from pre-trained models, and (2) selecting the state with the highest value according to the value function.
@@
To tackle these challenges, we propose a novel method, Soft Value-based Decoding in Diffusion models (SVDD), for optimizing downstream reward functions in diffusion models (Figure 1). Inspired by recent literature on RL-based fine-tuning (Uehara et al., 2024), we first introduce soft value functions that serve as look-ahead functions, indicating how intermediate noisy samples lead to high rewards in the future of the diffusion denoising process. After learning (or approximating) these value functions, we present a new inference-time technique, SVDD, which obtains multiple noisy states from the policy (i.e., denoising map) of pre-trained diffusion models and selects the sample with the highest value function at each time step. Specifically, we introduce two algorithms (SVDD-MC and SVDD-PM) depending on how we estimate value functions. Notably, the SVDD-PM approach does not require any additional learning as long as we have access to the reward feedback by utilizing the characteristics of diffusion models (i.e., the forward process in diffusion models to directly map $t$ to 00 in terms of expectation in Figure 1).
@@
Table 1: A comparison of SVDD to prior approaches. “Non-differentiable” refers to the method’s ability to operate without requiring differentiable proxy models. “No Training” means that no additional training of the diffusion model is required as long as we have access to the reward feedback. We compare to SMC-based methods in Section 6.
@@
Classifier guidance $\checkmark$
@@
Classifier-free $\checkmark$
@@
SVDD-MC (here) $\checkmark$ $\checkmark$
SVDD-PM (here) $\checkmark$ $\checkmark$ $\checkmark$
Our novel technique for optimizing downstream reward functions in pre-trained diffusion models makes two contributions (Table 1). First, it eliminates the need to construct differentiable proxy models. This allows for the use of non-differentiable reward feedback, which is common in many scientific fields, and makes our method applicable to recent discrete diffusion (Shi et al., 2024; Sahoo et al., 2024) models without any modification. Second, it avoids the need to fine-tune the generative model itself. This addresses the high computational cost associated with fine-tuning diffusion models. We demonstrate the effectiveness of our methods across diverse domains, including image generation, molecule generation (optimization of docking scores, QED, and SA), and DNA/RNA generation (optimization of activity levels).
@@ 2 Related Works
To summarize related work, we first outline methods relevant to our goal, categorizing them based on whether they involve fine-tuning. We then discuss related directions, such as discrete diffusion models, where our method excels. We defer other relevant works (e.g., decoding in LLMs) to Section A due to space constraints.
@@
We discuss two main methods for optimizing rewards in diffusion models without fine-tuning. We further cover closely relevant methods based on sequential Monte Carlo (SMC) in Section 6 and Appendix B, after presenting our method.
  • Classifier guidance (Dhariwal and Nichol, 2021; Song et al., 2020): It has been widely used to condition pre-trained diffusion models without fine-tuning. Although these methods do not originally focus on optimizing reward functions, they can be applied for this purpose (Uehara et al., 2024, Section 6.2). In this approach, an additional derivative of a certain value function is incorporated into the drift term (mean) of pre-trained diffusion models during inference. Subsequent variants (e.g., Chung et al. (2022); Ho et al. (2022); Bansal et al. (2023); Guo et al. (2024); Wang et al. (2022); Yu et al. (2023)) have been proposed to simplify the learning of value functions. However, these methods require constructing differentiable models, which limits their applicability to non-differentiable features/reward feedbacks commonly encountered in scientific domains as mentioned in Section 1. Additionally, this approach cannot be directly extended to discrete diffusion models (e.g., (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024)) in a principle way. Our approach aims to address these challenges.
    Note a notable exception of classifier guidance tailored to discrete diffusion models has been recently proposed by Nisonoff et al. (2024). However, SVDD can be applied to both continuous and discrete diffusion models in a unified manner. Furthermore, their practical method requires the differentiability of proxy models, unlike SVDD. We compare its performance with our method in Section 7. We provide further details in Section C.
  • Best-of-N: The naive way is to generate multiple samples and select the top samples based on the reward functions, known as best-of-N in the literature on (autoregressive) LLMs (Stiennon et al., 2020; Nakano et al., 2021; Touvron et al., 2023; Beirami et al., 2024; Gao et al., 2023). This approach is significantly less efficient than ours, as SVDD uses soft-value functions that predict how intermediate noisy samples lead to high rewards in the future. We validate this experimentally in Section 7.
@@
Several methods exist for fine-tuning generative models to optimize downstream reward functions, such as classifier-free guidance (Ho and Salimans, 2022) and RL-based fine-tuning (Fan et al., 2023; Black et al., 2023) or its variants (Dong et al., 2023; Wallace et al., 2024). However, these approaches often come with caveats, including high computational costs and the risk of easily forgetting pre-trained models. In our work, we propose an inference-time technique that eliminates the need for fine-tuning generative models.
@@
Based on seminal works Austin et al. (2021); Campbell et al. (2022), recent work on masked diffusion models (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024) has demonstrated their strong performance in natural language generation. Additionally, they have been applied to biological sequence generation (e.g., DNA, protein sequences in Campbell et al. (2024); Sarkar et al. (2024)). In these cases, the use of diffusion models over autoregressive models is particularly apt, given that many biological sequences ultimately adopt complex three-dimensional structures. We also note that ESM3 (Hayes et al., 2024), a widely recognized foundation model in protein sequence generation, bears similarities to masked diffusion models. Despite its significance, it cannot be integrated with standard classifier guidance, because adding a continuous gradient to a discrete objective is not inherently valid. Unlike standard classifier guidance, SVDD can be seamlessly applied to discrete diffusion models.
@@ 3 Preliminaries and Goal
We describe the standard method for training diffusion models and outline the objective of our work: optimizing downstream reward functions given pre-trained diffusion models.
@@ 3.1 Diffusion Models
In diffusion models (Ho et al., 2020; Song et al., 2020), our goal is to learn a sampler $p(x)\in\Delta(\mathcal{X})$ given data consisting of $x\in\mathcal{X}$. The training process for a standard diffusion model is summarized as follows. First, we introduce a (fixed) forward process $q_{t}:\mathcal{X}\to\Delta(\mathcal{X})$. Next, we aim to learn a backward process: $\{p_{t}\}$ where each $p_{t}$ is $\mathcal{X}\to\Delta(\mathcal{X})$ so that the distributions induced by the forward process and backward process match marginally. For this purpose, by parametrizing the backward processes with $\theta\in\mathbb{R}^{d}$, we typically use the following loss function:
@@
which is derived from the variational lower bound of the negative log-likelihood (i.e., ELBO).
@@
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. Here, $\hat{x}_{0}(x_{t};\theta)$ is a neural network that predicts $x_{0}$ from $x_{t}$ ($\mathbb{E}_{q}[x_{0}|x_{t}]$).
@@ 3.2 Objective: Generating Samples with High Rewards While Preserving Naturalness
We consider a scenario where we have a pre-trained diffusion model, which is trained using the loss function explained in Section 3.1. These pre-trained models are typically designed to excel at characterizing the natural design space (e.g., image space, biological sequence space, or chemical space) by emulating the extensive training dataset. Our work focuses on obtaining samples that also optimize downstream reward functions $r:\mathcal{X}\to\mathbb{R}$ (e.g., Quantitative Estimate of Druglikeness (QED) and Synthetic Accessibility (SA) in molecule generation), while maintaining the naturalness by leveraging pre-trained diffusion models. We formalize this goal as follows.
@@
$p^{(\alpha)}(x)$ $:=\mathop{\mathrm{argmax}}_{p\in[\Delta(\mathcal{X})]}\underbrace{\mathbb{E}_{x\sim p(\cdot)}[r(x)]}_{\textbf{term (a)}}-\alpha\underbrace{\mathrm{KL}(p(\cdot)\|p^{\mathrm{pre}}(\cdot))}_{\textbf{term(b)}}\propto\exp(r(x)/\alpha)p^{\mathrm{pre}}(x).$
Here, term (a) is introduced to optimize the reward function, while term (b) is used to maintain the naturalness of the generated samples.
@@
Several existing approaches target this goal (or its variant), as discussed in Section 2, including classifier guidance, fine-tuning (RL-based or classifier-free), and Best-of-N. In our work, we focus on non-fine-tuning-based methods; specifically, we aim to address the limitations of these methods: the requirement for differentiable proxy models in classifier guidance and the inefficiency of Best-of-N.
Finally, we note that all results discussed in this paper can be easily extended to cases where the pre-trained model is a conditional diffusion model. For example, in our image experiments (Section 7), the pre-trained model is a conditional diffusion model conditioned on text (e.g., Stable Diffusion).